Given that C= 500+ 0.8y, if the level of disposable income is 1000, The level of saving is?
1
Expert's answer
2012-03-16T09:03:15-0400
Let’s indicate, that Y is a level of disposable income, S is the level of saving and C is the level of consumption, then:
Y = S + C
S = Y – C
For the disposable income (1000) the level of savings is:
S = y – C = y – (500+ 0.8y) = 0,2y – 500 = 0,2 * 1000 – 500 = 200 – 500 = -300.
The level of savings is -300, this means that people live in debt.